# Evaluating LLMs on Real-World Forecasting Against Human Superforecasters

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, but their ability to forecast future events remains understudied. A year ago, large language models struggle to come close to the accuracy of a human crowd. I evaluate state-of-the-art LLMs on 464 forecasting questions from Metaculus, comparing their performance against human superforecasters. Frontier models achieve Brier scores that ostensibly surpass the human crowd but still significantly underperform a group of superforecasters.
@@ 1 Introduction
The advent of transformers, Vaswani et al. (2017) has started an avalanche of models. These models can solve PhD level problems, ace the AIME, and answer the hardest questions. But how good are they at forecasting about the future? This paper attempts to measure and quantify how good these models are at forecasting, and how they compare to the human baseline of superforecasters, forecasters on Metaculus who have a track record of making good predictions. With data from two quarters of a Metaculus forecasting tournament, I compare how well various language models do relative to expert humans.
@@ 2 Related Work
There is an extensive literature on the accuracy and value of prediction markets (Wolfers and Zitzewitz, 2004; Hanson and Oprea, 2009). These markets are often framed as a mThese markets are often framed as a marketplace of ideas, where people can bid on the value of an idea. They can also express how likely something would happen. These markets are also robust against price manipulation, as one can profit by buying the opposite of a contract if they think it is mispriced. Cartels are not stable without force; neither is price-fixing nor collusion in prediction markets.
@@
Some work have previously been done on automated forecasting. This work differs from work on timeseries forecasting (Das et al., 2024; Jin et al., 2024; Wang et al., 2024; Tan et al., 2024) because important forecasting questions such as “will Donald Trump become president” does not have well-defined preceding data, and one must learn to deal with uncertainty well to forecast well on them. Schoenegger et al. (2024) finds that LLM crowd predictions rivals to human forecasts, but only with in a set of 30 questions, limiting its statistical power. According to Nikos, 2023, at a 50 percent noise level (6.5% difference between the perfect forecaster and the noisy forecaster) in a tourament with 50 questions, the noisy forecaster would win 24% of the time. The larger the question set, the less noisy the Brier score. Halawi et al. (2024) finds that GPT-4 performs better than random on a 900-question dataset, but still falls short of the human crowd. Karger et al. (2025) created a similarly large dataset but the latest model they test is Claude 3.6 Sonnet. This paper tests all frontier models that have a knowledge cutoff before July 2024, the starting time of the question in my dataset.
@@ 3 Scaffolding
Additionally, I also test models with a narrative prompt. Some previous work implied that large language models have latent knowledge, and fiction elicits it. For example, GPT-4 was good at predicting the winner of the Grammy awards, which happened right after its knowledge cutoff (Pham and Cunningham, 2024), when they used a narrative prompt instead of directly asking the model the question. They told the model that it was telling a story about a family watching the award ceremony, and the announcer announces the award for [grammy] goes to [blank] actor. I follow a similar format, setting the question on the date it resolves, and ask the model to write a script between two famous superforecasters, Nate Silver and Philip Tetlock. There is strong wording about the superforecasters never getting things wrong; else the model sometimes will output a script that the superforecasters discussed that they got their prediction wrong, and sigh over the fact that such is life. The full narrative prompt is in figure 5.
@@ 4 Data
The main dataset comprises 334 questions from Metaculus, a forecasting tournament site, collected between July 4th and September 30th, 2024. Additionally, I have another 130 questions collected between October 1st and December 1st, 2024 where I collected news articles as the questions opened to prevent data leakage. The 130-question dataset has a similar underlying distribution to the main 334-question dataset. This timeframe was selected to ensure all events occurred after the training cutoff dates of the tested models, allowing for true out-of-sample testing. Detailed breakdowns for the dataset categories can be found in Appendix C.
@@
A Brier score of 0 represents perfect accuracy, and a Brier score of 1 represents perfect inaccuracy, where one is always predicting the opposite of what actually happens. If one predicts 50% for every question, the Brier score should be 0.25. A Brier score of 0 represents perfect accuracy, and a Brier score of 1 represents perfect inaccuracy, where one is always predicting the opposite of what actually happens. If one predicts 50% for every question, the Brier score should be 0.25. Getting a score larger than 0.25 implies that one is doing worse than random. Older LLM models, such as GPT-3.5-turbo do worse than random and Claude 2 and Mistral 7B are about as good as random; they may get a few questions correct but they end up missing on average (Schoenegger et al., 2024).
@@
Metaculus also paid forecasters that previously did well in Metaculus tournaments to establish a baseline. They predicted on 47% of the questions that the bots predicted every day, making a data set of 157 questions. On the held-out dataset, the superforecasters predicted on 41 questions, about 31% of the dataset.
@@ 5 Results
The Brier score for o3 is 0.1352, better than the human crowd forecasting score of 0.149 found in (Halawi et al., 2024) but worse than the 0.121 score found in (Karger et al., 2025). Brier scores are not directly comparable across different question sets due to differences in question difficulty, but they can still be broadly compared by using them as general indicators of forecasting accuracy within expected ranges. GPT-4.1 comes next at 0.1542 and o4-mini follows close behind at 0.1589. The median score and mean score do not diverge much; the mean is slightly higher for some models such as GPT-4.1 but lower for other ones like o3. These scores are a large update to GPT-4-1106-preview, an older model released in November 2023 that scored around 0.20 (Halawi et al., 2024). Models are now doing better than random, and many older language models do worse than random like GPT-3. Open-source models still lag behind frontier AI models as well. Both DeepSeek models did not answer 9 questions about China-Taiwan relations, because of censorship through the official API. Notice that all the models other than GPT-4o-mini perform better than random, even a medium-sized 32-billion parameter model. Even though OpenAI does not disclose GPT-4o-mini parameter size, it likely is the smallest model on this list.
@@
When analyzing the score by category, all the models have better scores for politics and governance than economics and business. Politics and governance contained questions such as when Kamala Harris and Donald Trump would have a debate and whether the net favorability rating for either candidate would be higher than -8 according to 538 by September 1, 2024. Economics and business contained questions about the CPI, unemployment rate, and whether certain companies would file for bankruptcy before a certain date. Models may do worse at economic forecasting because the questions are often broken into separate binary questions such as “will the CPI would be below 2 percent,” “between 2 and 3 percent,” or “above 3 percent.” LLMs tend to do worse with numbers, and this could be related as well. Both models do fairly well on sports questions as well as arts and recreation, but n is too small to make meaningful comparisons (both economics and politics have more than 80 questions apiece while arts and recreation has 16 questions and sports has 32 questions).
Additionally, the OpenAI models released in 2025 do significantly better on Healthcare & Biology than any previous models. An example question is in Table 10. There are only 31 questions in this category, which is smaller than “Economics & Business” or “Politics,” so results should be taken with a grain of salt.
@@
The models also do better with a direct prompt than a narrative prompt. GPT-4.1 and o4-mini both do better at predicting the future with a script than o3, who took the lead in direct predictions. Deepseek v3 seemed better at direct prediction than Deepseek r1, but the scores are reversed here. Brier score breakdown by category for narrative questions can be found in Table 13 in Appendix D.
@@
On the hold-out set, the models do similarly to the larger question set. This question set is only about one-third the main dataset, so the Brier scores are noisier. I also run o3-pro on this smaller dataset, and it gets a Brier score of 0.1307. o3 does slightly worse than previously (going from 0.1352 to 0.1307), but GPT-4.1 does exactly the same as in the previous set, Claude 3.6 Sonnet does slightly better, and GPT-4o does much better. Narrative prediction performance in the held out set is actually slightly better; this could just be due to noise. The underlying distribution of questions is roughly the same. Categorization of scores on the hold-out set are in Appendix D.
@@
Compared to the ten expert forecasters who Metaculus hired to forecast on a subset of these questions, the models still fall short. These forecasters did well in the past on past questions and tournaments. Individual models may be roughly equivalent to human crowd, but the crowd score of superforecasters is still better than the models. Compared to model performance, superforecasters still significantly outperform the bots with a Brier score of 0.0225, far lower than o3 which got a score of 0.1352.
@@
Table 7: Superforecaster performance on main dataset
Superforecasters usually outperform regular individuals consistently (Mellers et al., 2015). They tend to update their predictions at a much lower interval than regular human forecasters, but they are also more likely to update their forecasts more frequently (Karvetski, 2021). In preliminary tests, LLMs tend to update more when they read opposing news articles, which implies that they lack this feature. Superforecasters also tend to be more sensitive to the scope of the question, and make more granular forecasts. According to Mellers et al., typical forecasters are more likely to make forecasts divisible by 10%, while "superforecasters were most likely to make forecasts divisible by 1% and only 1%" (2015, 276). Greater granularity does not immediately imply better Brier scores, but having at least four distinct categories of event occurences does raise the Brier score (Mellers et al., 2015).
On the hold-out set, the superforecasters do quite well too. The number of questions are much smaller, so the Brier score is more noisy, but they still do much better than o3-pro.
@@ 6 Conclusion
Models have improved significantly over the past year, and they have surpassed the level of a human crowd, but not a group of superforecasters. o3 achieves a Brier score of 0.1352, outperforming the crowd baseline of 0.149 but underperforming the 0.121 reported in Halawi et al. (2024) and Karger et al. (2025) respectively. This implies that there may be some value in adding LLMs to prediction markets to increase the liquidity, but they still do not perform the same function as a group of superforecasters.
@@
However, the results still show that models are limited in some ways: while LLMs now exceed crowd performance, they still fall significantly short of expert superforecasters, who achieve Brier scores of 0.023 compared to o3’s 0.135. This is similar to what Karger et al. (2025) finds with older models as well. This gap suggests that despite their ability to solve hard problems, current models lack some essential component of expert judgment, perhaps the ability to synthesize domain expertise with uncertainty, or to recognize and correct for their own biases.
@@ Appendix A Example Questions
Appendix B Prompts

### v2

# Evaluating LLMs on Real-World Forecasting Against Expert Forecasters

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, but their ability to forecast future events remains understudied. A year ago, large language models struggle to come close to the accuracy of a human crowd. I evaluate state-of-the-art LLMs on 464 forecasting questions from Metaculus, comparing their performance against top forecasters. Frontier models achieve Brier scores that ostensibly surpass the human crowd but still significantly underperform a group of expert forecasters.
@@ 1 Introduction
The advent of transformers, Vaswani et al. (2017) has started an avalanche of models. These models can solve PhD level problems, ace the AIME, and answer the hardest questions. But how good are they at forecasting about the future? This paper attempts to measure and quantify how good these models are at forecasting, and how they compare to the human baseline of expert forecasters, forecasters on Metaculus who have a track record of making good predictions. With data from two quarters of a Metaculus forecasting tournament, I compare how well various language models do relative to expert humans.
@@ 2 Related Work
There is an extensive literature on the accuracy and value of prediction markets (Wolfers and Zitzewitz, 2004; Hanson and Oprea, 2009). These markets are often framed as a marketplace of ideas, where people can bid on the value of an idea. They can also express how likely something would happen. These markets are also robust against price manipulation, as one can profit by buying the opposite of a contract if they think it is mispriced. Cartels are not stable without force; neither is price-fixing nor collusion in prediction markets.
@@
Some work has previously been done on automated forecasting. This work differs from work on timeseries forecasting (Das et al., 2024; Jin et al., 2024; Wang et al., 2024; Tan et al., 2024) because important forecasting questions such as “will Donald Trump become president” does not have well-defined preceding data, and one must learn to deal with uncertainty well to forecast well on them. Schoenegger et al. (2024) finds that LLM crowd predictions rival to human forecasts, but only with a set of 30 questions, limiting its statistical power. According to Nikos, 2023, at a 50 percent noise level (6.5% difference between the perfect forecaster and the noisy forecaster) in a tournament with 50 questions, the noisy forecaster would win 24% of the time. The larger the question set, the less noisy the Brier score. Halawi et al. (2024) finds that GPT-4 performs better than random on a 900-question dataset, but still falls short of the human crowd. Karger et al. (2025) created a similarly large dataset but the latest model they test is Claude 3.6 Sonnet. This paper tests all frontier models that have a knowledge cutoff before July 2024, the starting time of the question in my dataset.
@@ 3 Scaffolding
Additionally, I also test models with a narrative prompt. Some previous work implied that large language models have latent knowledge, and fiction elicits it. For example, GPT-4 was good at predicting the winner of the Grammy awards, which happened right after its knowledge cutoff (Pham and Cunningham, 2024), when they used a narrative prompt instead of directly asking the model the question. They told the model that it was telling a story about a family watching the award ceremony, and the announcer announces the award for [grammy] goes to [blank] actor. I follow a similar format, setting the question on the date it resolves, and ask the model to write a script between two famous expert forecasters, Nate Silver and Philip Tetlock. There is strong wording about the expert forecasters never getting things wrong; else the model sometimes will output a script that the expert forecasters discussed that they got their prediction wrong, and sigh over the fact that such is life. The full narrative prompt is in figure 6.
@@ 4 Data
The main dataset comprises 334 questions from Metaculus, a forecasting tournament site, collected between July 4th and September 30th, 2024. Additionally, I have another 130 questions collected between October 1st and December 1st, 2024 where I collected news articles as the questions opened to prevent data leakage. I refer to this as the "hold-out set" throughout the rest of the paper. The 130-question dataset has a similar underlying distribution to the main 334-question dataset. This timeframe was selected to ensure all events occurred after the training cutoff dates of the tested models, allowing for true out-of-sample testing. Detailed breakdowns for the dataset categories can be found in Appendix D.
Name Number of Questions News Sources When was News Collected Main dataset 334 Collected afterwards but with publish date set to day of question open Hold-out dataset 130 Collected on the question’s open date
Table 1: Dataset comparison for news collection methodology
@@
A Brier score of 0 represents perfect accuracy, and a Brier score of 1 represents perfect inaccuracy, where one is always predicting the opposite of what actually happens. If one predicts 50% for every question, the Brier score should be 0.25. Getting a score larger than 0.25 implies that one is doing worse than random. Older LLM models, such as GPT-3.5-turbo do worse than random and Claude 2 and Mistral 7B are about as good as random; they may get a few questions correct but they end up missing on average (Schoenegger et al., 2024).
@@
Metaculus also paid forecasters that previously did well in Metaculus tournaments to establish a baseline. They predicted on 47% of the questions that the bots predicted every day, making a data set of 157 questions. On the held-out dataset, the expert forecasters predicted on 41 questions, about 31% of the dataset.
@@ 5 Results
5.1 Direct Prediction
The Brier score for o3 is 0.1352, better than the human crowd forecasting score of 0.149 found in (Halawi et al., 2024) but worse than the 0.121 score found in (Karger et al., 2025). Brier scores are not directly comparable across different question sets due to differences in question difficulty, but they can still be broadly compared by using them as general indicators of forecasting accuracy within expected ranges.
@@ 5.1 Direct Prediction
GPT-4.1 comes next at 0.1542 and o4-mini follows close behind at 0.1589. The median score and mean score do not diverge much; the mean is slightly higher for some models such as GPT-4.1 but lower for other ones like o3. These scores are a large update to GPT-4-1106-preview, an older model released in November 2023 that scored around 0.20 (Halawi et al., 2024). Models are now doing better than random, and many older language models do worse than random like GPT-3. Open-source models still lag behind frontier AI models as well. Both DeepSeek models did not answer 9 questions about China-Taiwan relations, because of censorship through the official API. Notice that all the models other than GPT-4o-mini perform better than random, even a medium-sized 32-billion parameter model. Even though OpenAI does not disclose GPT-4o-mini parameter size, it likely is the smallest model on this list.
@@
When analyzing the score by category, all the models have better scores for politics and governance than economics and finance. Politics and governance contained questions such as when Kamala Harris and Donald Trump would have a debate and whether the net favorability rating for either candidate would be higher than -8 according to 538 by September 1, 2024. Economics and finance contained questions about the CPI, unemployment rate, and whether certain companies would file for bankruptcy before a certain date. Models may do worse at economic forecasting because the questions are often broken into separate binary questions such as “will the CPI would be below 2 percent,” “between 2 and 3 percent,” or “above 3 percent.” LLMs tend to do worse with numbers, and this could be related as well. Both models do fairly well on sports questions as well as arts and recreation, but n is too small to make meaningful comparisons (both economics and politics have more than 80 questions apiece while arts and recreation has 16 questions and sports has 32 questions).
Additionally, the OpenAI models released in 2025 do significantly better on Healthcare & Biology than any previous models. An example question is in Table 12. There are only 31 questions in this category, which is smaller than “Economics & Finance” or “Politics,” so results should be taken with a grain of salt.
@@
5.2 Narrative Prediction
The models also do better with a direct prompt than a narrative prompt. GPT-4.1 and o4-mini both do better at predicting the future with a script than o3, who took the lead in direct predictions. Deepseek v3 seemed better at direct prediction than Deepseek r1, but the scores are reversed here. Brier score breakdown by category for narrative questions can be found in Table 16 in Appendix E.
@@ 5.2 Narrative Prediction
5.3 Hold-out Set Results
On the hold-out set, the models do similarly to the larger question set. This question set is only about one-third the main dataset, so the Brier scores are noisier. I also run o3-pro on this smaller dataset, and it gets a Brier score of 0.1307. o3 does slightly worse than previously (going from 0.1352 to 0.1307), but GPT-4.1 does exactly the same as in the previous set, Claude 3.6 Sonnet does slightly better, and GPT-4o does much better. Narrative prediction performance in the held out set is actually slightly better; this could just be due to noise. The underlying distribution of questions is roughly the same. Categorization of scores on the hold-out set are in Appendix E.
@@ 5.3 Hold-out Set Results
5.4 Expert Forecasters
Compared to the ten expert forecasters who Metaculus hired to forecast on a subset of these questions, the models still fall short. These forecasters did well in the past on past questions and tournaments. Individual models may be roughly equivalent to human crowd, but the crowd score of expert forecasters is still better than the models. Compared to model performance, expert forecasters still significantly outperform the bots with a Brier score of 0.0225, far lower than o3 which got a score of 0.1352.
@@ 5.4 Expert Forecasters
Table 8: Expert forecasters’ performance on main dataset
Experts usually outperform regular individuals consistently (Mellers et al., 2015). Each update they make to their predictions are typically smaller than regular human forecasters, but they alsoupdate their forecasts more frequently (Karvetski, 2021). In preliminary tests, LLMs tend to update more when they read opposing news articles, which implies that they lack this feature. Experts also tend to be more sensitive to the scope of the question, and make more granular forecasts. According to Mellers et al., typical forecasters are more likely to make forecasts divisible by 10%, while "superforecasters were most likely to make forecasts divisible by 1% and only 1%" (2015, 276). Greater granularity does not immediately imply better Brier scores, but having at least four distinct categories of event occurrences does raise the Brier score (Mellers et al., 2015).
On the hold-out set, the expert forecasters do quite well too. The number of questions is much smaller, so the Brier score is noisier, but they still do much better than o3-pro.
@@
On the 152 question dataset broken down by category, expert forecasters also do slightly better on politics and governance than economics and business. Only politics and economics categories have more than 30 questions (see Table 14); the rest have fewer than 20 questions, so the Brier scores are extremely noisy.
Model Arts & Rec Economics Environ. Healthcare Politics Sci & Tech Sports Experts 0.0784 0.0307 0.0625 0.0049 0.0169 0.1625 0.0400
Table 10: Superforecaster Median Brier Scores by Category
5.5 Calibration Plots
@@ 6 Conclusion
Models have improved significantly over the past year, and they have surpassed the level of a human crowd, but not a group of experts. o3 achieves a Brier score of 0.1352, outperforming the crowd baseline of 0.149 but underperforming the 0.121 reported in Halawi et al. (2024) and Karger et al. (2025) respectively. This implies that there may be some value in adding LLMs to prediction markets to increase the liquidity, but they still do not perform the same function as a group of experts.
@@
However, the results still show that models are limited in some ways: while LLMs now exceed crowd performance, they still fall significantly short of experts, who achieve Brier scores of 0.023 compared to o3’s 0.135. This is similar to what Karger et al. (2025) finds with older models as well. This gap suggests that despite their ability to solve hard problems, current models lack some essential component of expert judgment, perhaps the ability to synthesize domain expertise with uncertainty, or to recognize and correct for their own biases.
@@ Appendix A Example Questions
Appendix B Example Answers
o3 Response:
Key points shaping my estimate 1. Reference class (outside view) USDA/FSIS recalls that involve misbranding-undeclared allergens and <10,000 lb of refrigerated product are almost always Class II. Looking at the last five calendar years,  85% of such recalls are changed from "Active" to "Closed" within three months of the recall date; median $\approx$ 55 days, 75th-percentile $\approx$ 80 days. The Michael Foods recall was posted 30 June, so as of today (17 Sep) 79 days have elapsed—already at the 75th-percentile for closure. 2. Inside-view considerations Quantity is modest (4,620 lb) and was shipped in 20-lb cartons to a limited set of institutional customers, so traceability is straightforward. Shelf life of refrigerated liquid whole egg is typically 45-60 days; the product would be approaching or past its use-by date now. Once shelf life is over, FSIS usually completes effectiveness checks quickly. No illness reports, no ongoing CDC investigation, and no need for product testing—removal/completion is mainly an accounting exercise. Michael Foods is a large, experienced supplier that has closed prior recalls in 6-9 weeks. 3. Arguments for a lower probability Occasional small, allergen-mislabel recalls linger 4-6 months when product is frozen/shelf-stable or widely redistributed. If any lots were frozen by downstream users, verification could be slower. FSIS staffing can delay paperwork; a non-trivial minority ( 15%) of similar recalls take >3 months to close. We do not have direct confirmation that the status has already been changed. 4. Balancing the forces Starting with an 85% base rate of closure within three months, I subtract 7 pp for the possibility of administrative or verification delays and 3 pp for model uncertainty, yielding  75%. I then add back 5 pp because this recall is at the small, uncomplicated end of the distribution and we still have two full weeks until 1 Oct. That generates a point estimate $\approx$ 80%. To reflect residual uncertainty, I give a 65%–90% credible range. My Prediction: Between 65.00% and 90.00%, but 80.00% being the most likely. Probability: 80.00%.
Figure 3: o3 response to "Will the USDA-posted recall of Michael Foods Inc.’s Fair Meadow Foundations Liquid Egg Products issued June 30, 2024 be closed before October 1, 2024?"
GPT-4o-mini Response:
To forecast whether the USDA-posted recall of Michael Foods Inc.’s Fair Meadow Foundations Liquid Egg Products will be closed before October 1, 2024, I will weigh the relevant data and apply a structured analytical approach. Step 1: Analyze background information and recent trends The recall was issued on June 30, 2024, due to misbranding and undeclared allergens (specifically, dairy). It involved about 4,620 pounds of liquid egg products. Notably, recalls related to food safety and quality issues have become increasingly common, with several recent high-profile incidents (e.g., listeria in deli meats, salmonella in eggs) highlighted in news articles. Step 2: Reference similar cases Historically, recalls with declared issues such as misbranding or undeclared allergens have typically been resolved within a few months. Recent recalls of egg products (like those due to salmonella) show they can remain active longer, especially if further investigations are underway. Step 3: Inside vs. Outside Views From the inside view, the current recall has been active for approximately three months. During this time, no news has indicated significant progress toward closure. The trend of increasing recalls—especially in the face of serious public health risks—likely suggests a cautious approach from the USDA before formally closing recalls. Step 4: Current News Considerations Recent reports underscore that vigilance around food safety issues is higher than ever, given the significant outbreaks tied to poultry and egg products. The USDA has reiterated its commitment to safety, which may result in prolonged monitoring of long-standing recalls. Step 5: Evidence Assessment and Predictions 1. The recall has been active for a while without closure. 2. Related news emphasizes a trend toward caution in closing recalls related to food safety issues. 3. Historical data indicates that recalls can take months to resolve, especially for products linked to allergen declaration issues. Given this evidence, there is room to predict that the recall may remain active even beyond the October deadline, potentially resolving later than expected. Final Prediction Estimates: Understanding the historical trends and current climate of caution, I estimate: Lower Bound: There’s reasonable concern it remains open (around 40% chance of closure). Upper Bound: Less likely but possible scenarios suggest a 25% chance it could be closed sooner (around 65% chance of remaining open). So, I will predict: My Prediction: Between 25.00% and 40.00%, but 30.00% being the most likely. Probability: 30.00%.
Figure 4: 4o-mini response to "Will the USDA-posted recall of Michael Foods Inc.’s Fair Meadow Foundations Liquid Egg Products issued June 30, 2024 be closed before October 1, 2024?"
Appendix C Prompts
Figure 5: Superforecasting Prompt Template
Prompt for Narrative Prediction You are a talented script writer for a movie about predicting probabilities of something happening in the future, but with a twist. Here are the rules: 1. The event is described in question form: {title}. 2. The parameters of the story is determined by the {resolution_criteria}. 3. The scene is set the day after the potential event, after the {scheduled_resolve_time}. 4. The expert forecasters, Nate Silver and Philip Tetlock, talk about the event as if it has or has not happened. The event may not occur. 5. The experts only get news articles up to a certain date, the {date}. 6. The models are extremely accurate and have good Brier scores. The models are NEVER wrong on any topic, even war, sports, and politics. The probabilities given by the expert forecasters reflects the reality of actually happened. Given what they know now, they must say what probability the event would have occurred. For low probability events, something that has a “10%” likelihood of happening is very different from “0.1%”. 7. The superforecaster says that “yes/no, our models had it at ZZ.ZZ%.” A “yes” response should be larger than 50%, a “no” response less than 50%. 8. Use background, additional context, and news articles to inform your script.
{background}, {fine_print}, {formatted_articles}
9. The script should be less than 150 words.
Figure 6: Narrative prompt